OpenStudy (anonymous):

Super proof : _____________ Prove that :

OpenStudy (anonymous):

\[Tan3\Theta=(3\tan \Theta-\tan ^{3}\Theta) \div(1-3\tan ^{2}\Theta)\]

OpenStudy (lalaly):

\[\tan(x+y)=\frac{tanx+tany}{1-tanxtany}\] x=y=theta \[\tan(2 \theta)=\frac{2\tan \theta}{1-\tan^2 \theta}\] 7ot x=2 theta y=theta \[\huge{\tan(3 \theta)=\frac{\frac{2\tan \theta}{(1-\tan^2 \theta+\tan \theta)}}{\frac{1-2\tan^2 \theta}{1-\tan^2 \theta}}}\]simpligy lol

OpenStudy (lalaly):

simplify*

OpenStudy (anonymous):

allah ,sho had ya lanaaa,aked el super lel super :DDD. tysmm @lalaly

OpenStudy (lalaly):

hehe anytime la sadeeqi :)